# MaIL: Improving Imitation Learning with Mamba

### v1

@@ Abstract
This work introduces Mamba Imitation Learning (MaIL), a novel imitation learning (IL) architecture that offers a computationally efficient alternative to state-of-the-art (SoTA) Transformer policies. Transformer-based policies have achieved remarkable results due to their ability in handling human-recorded data with inherently non-Markovian behavior. However, their high performance comes with the drawback of large models that complicate effective training. While state space models (SSMs) have been known for their efficiency, they were not able to match the performance of Transformers. Mamba significantly improves the performance of SSMs and rivals against Transformers, positioning it as an appealing alternative for IL policies. MaIL leverages Mamba as a backbone and introduces a formalism that allows using Mamba in the encoder-decoder structure. This formalism makes it a versatile architecture that can be used as a standalone policy or as part of a more advanced architecture, such as a diffuser in the diffusion process. Extensive evaluations on the LIBERO IL benchmark and three real robot experiments show that MaIL: i) outperforms Transformers in all LIBERO tasks, ii) achieves good performance even with small datasets, iii) is able to effectively process multi-modal sensory inputs, iv) is more robust to input noise compared to Transformers.
@@ 3.1 Mamba: Selective State-Space Models
Inspired by the Attention mechanism [15] in Transformers, Mamba [13] improves upon the Structured State-Space Sequence Models (SSMs) [12] by using a selective scan operator, to propagate or forget information over time, allowing it to filter out relevant features. Specifically, we denote the input sequence $x\in\mathbb{R^{\mathtt{B}\times\mathtt{L}\times\mathtt{D}}}$ and the output sequence $y\in\mathbb{R^{\mathtt{B}\times\mathtt{L}\times\mathtt{D}}}$, where $\mathtt{B}$, $\mathtt{L}$, $\mathtt{D}$ refer to the batch size, sequence length and dimension respectively. A standard SSM with hidden state dimension $\mathtt{N}$ defines time-invariant parameters $\mathbf{A}\in\mathbb{R}^{\mathtt{N}\times\mathtt{N}}$, $\mathbf{B}\in\mathbb{R}^{\mathtt{N}\times\mathtt{D}}$, $\mathbf{C}\in\mathbb{R}^{\mathtt{D}\times\mathtt{N}}$ and the time-step vector $\mathbf{\Delta}$ to map the inputs up to $x(l)$ to a hidden state, which then can be projected to the output $y(l)$ (More details are in Appendix A). However, this property limits the SSM for contextual learning. Mamba implements the selection mechanism by making the SSM parameters a function of the input, that is,
@@ 5.2 Simulation Evaluation
LIBERO [14]: The evaluation is conducted using the LIBERO benchmark, which encompasses five distinct task suites: LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, LIBERO-Long, and LIBERO-90. Each task suite comprises 10 tasks along with 50 human demonstrations except for LIBERO-90 which contains 90 tasks with 50 demonstrations. Each task suite is designed to test different aspects of robotic learning and manipulation capabilities. The task visualizations are in Figure 3. More details can be found in Appendix C.
@@ 5.3 Real Robot Evaluation
ED-Tr $0.45$ $0.25$ $0.35$ $\mathbf{0.2}$
ED-Ma ss $\mathbf{0.55}$ $\mathbf{0.70}$ $\mathbf{0.45}$ $0.1$
@@ 7 Conclusion
In conclusion, this work presents MaIL, a novel imitation learning (IL) policy architecture that bridges the gap between efficiency and performance in handling sequences of observations. By leveraging the strengths of state space models and rigorously improving upon them, MaIL offers a competitive alternative to the traditionally large and complex Transformer-based policies. The introduction of Mamba in an encoder-decoder structure enhances its versatility, making it suitable for standalone use as well as integration into advanced architectures like diffusion processes. Extensive evaluations on the LIBERO IL benchmark and real robot experiments demonstrate that MaIL not only matches but also surpasses the performance of existing baselines, establishing it as a promising approach for IL tasks.
Appendix A State Space Models
@@ Appendix B Transformer Architecture
Appendix C Additional Task Details
C.1 LIBERO benchmark
C.2 Real Robot Tasks
C.2.1 CupStacking
C.2.2 Inserting
C.2.3 PickPlacing
@@ E.3 Data Collection
Figure 12: Data collection scene for real robot

### v2

@@ Abstract
This work presents Mamba Imitation Learning (MaIL), a novel imitation learning (IL) architecture that provides an alternative to state-of-the-art (SoTA) Transformer-based policies. MaIL leverages Mamba, a state-space model designed to selectively focus on key features of the data. While Transformers are highly effective in data-rich environments due to their dense attention mechanisms, they can struggle with smaller datasets, often leading to overfitting or suboptimal representation learning. In contrast, Mamba’s architecture enhances representation learning efficiency by focusing on key features and reducing model complexity. This approach mitigates overfitting and enhances generalization, even when working with limited data. Extensive evaluations on the LIBERO benchmark demonstrate that MaIL consistently outperforms Transformers on all LIBERO tasks with limited data and matches their performance when the full dataset is available. Additionally, MaIL’s effectiveness is validated through its superior performance in three real robot experiments. Our code is available at https://github.com/ALRhub/MaIL.
@@ 3.1 Mamba: Selective State-Space Models
Inspired by the Attention mechanism [15] in Transformers, Mamba [13] improves upon the Structured State-Space Sequence Models (SSMs) [12] by using a selective scan operator, to propagate or forget information over time, allowing it to filter out relevant features. Specifically, we denote the input sequence $x\in\mathbb{R^{\mathtt{B}\times\mathtt{L}\times\mathtt{D}}}$ and the output sequence $y\in\mathbb{R^{\mathtt{B}\times\mathtt{L}\times\mathtt{D}}}$, where $\mathtt{B}$, $\mathtt{L}$, $\mathtt{D}$ refer to the batch size, sequence length and dimension respectively. A standard SSM with hidden state dimension $\mathtt{N}$ defines time-invariant parameters $\mathbf{A}\in\mathbb{R}^{\mathtt{N}\times\mathtt{N}}$, $\mathbf{B}\in\mathbb{R}^{\mathtt{N}\times\mathtt{D}}$, $\mathbf{C}\in\mathbb{R}^{\mathtt{D}\times\mathtt{N}}$ and the time-step vector $\mathbf{\Delta}$ to map the inputs up to $x(l)$ to a hidden state, which then can be projected to the output $y(l)$ (More details are in Appendix C.1). However, this property limits the SSM for contextual learning. Mamba implements the selection mechanism by making the SSM parameters a function of the input, that is,
@@ 5.2 Simulation Evaluation
LIBERO [14]: The evaluation is conducted using the LIBERO benchmark, which encompasses five distinct task suites: LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, LIBERO-Long, and LIBERO-90. Each task suite comprises 10 tasks along with 50 human demonstrations except for LIBERO-90 which contains 90 tasks with 50 demonstrations. Each task suite is designed to test different aspects of robotic learning and manipulation capabilities. The task visualizations are in Figure 3. More details can be found in Appendix B.
@@ 5.3 Real Robot Evaluation
ED-Tr $0.45$ $0.25$ $\mathbf{0.70}$ $0.45$
ED-Ma ss $\mathbf{0.55}$ $\mathbf{0.70}$ $\mathbf{0.70}$ $\mathbf{0.55}$
@@ 7 Conclusion
In conclusion, this work presents MaIL, a novel imitation learning (IL) policy architecture that bridges the gap between efficiency and performance in handling sequences of observations. By leveraging the strengths of state space models and rigorously improving upon them, MaIL offers a competitive alternative to the traditionally large and complex Transformer-based policies. The introduction of Mamba in an encoder-decoder structure enhances its versatility, making it suitable for standalone use as well as integration into advanced architectures like diffusion processes. Extensive evaluations on the LIBERO benchmark and real robot experiments demonstrate that MaIL not only matches but also surpasses the performance of existing baselines with limited demonstrations, establishing it as a promising approach for IL tasks.
Appendix A More Experimental Results
A.1 Comparison on Different Checkpoints
Evaluating multiple checkpoints can provide important insights into our model’s generalization capabilities. Here, we conducted additional experiments assessing the model after 40, 50, and 60 epochs. The results presented in Table 5 suggest that the model neither overfits nor underfits the data at these intervals.
w/o language w/ language
DDP-H5 Backbone LIBERO-Object LIBERO-Spatial LIBERO-Long LIBERO-Goal
Epoch 40 ED-Tr $0.770\scriptstyle{\pm 0.063}$ $0.415\scriptstyle{\pm 0.085}$ $0.345\scriptstyle{\pm 0.044}$ $0.502\scriptstyle{\pm 0.074}$
ED-Ma $0.802\scriptstyle{\pm 0.030}$ $0.478\scriptstyle{\pm 0.039}$ $0.355\scriptstyle{\pm 0.026}$ $\mathbf{0.615\scriptstyle{\pm 0.028}}$
Epoch 50 ED-Tr $0.758\scriptstyle{\pm 0.045}$ $0.412\scriptstyle{\pm 0.060}$ $0.333\scriptstyle{\pm 0.023}$ $0.497\scriptstyle{\pm 0.072}$
ED-Ma $\mathbf{0.807\scriptstyle{\pm 0.051}}$ $\mathbf{0.500\scriptstyle{\pm 0.031}}$ $0.385\scriptstyle{\pm 0.023}$ $0.580\scriptstyle{\pm 0.041}$
Epoch 60 ED-Tr $0.767\scriptstyle{\pm 0.045}$ $0.448\scriptstyle{\pm 0.092}$ $0.348\scriptstyle{\pm 0.050}$ $0.505\scriptstyle{\pm 0.036}$
ED-Ma $0.805\scriptstyle{\pm 0.058}$ $0.497\scriptstyle{\pm 0.050}$ $\mathbf{0.390\scriptstyle{\pm 0.028}}$ $0.610\scriptstyle{\pm 0.035}$
Max ED-Tr $0.770\scriptstyle{\pm 0.063}$ $0.448\scriptstyle{\pm 0.092}$ $0.348\scriptstyle{\pm 0.05}$ $0.505\scriptstyle{\pm 0.036}$
ED-Ma $\mathbf{0.807\scriptstyle{\pm 0.051}}$ $\mathbf{0.500\scriptstyle{\pm 0.031}}$ $\mathbf{0.390\scriptstyle{\pm 0.028}}$ $\mathbf{0.615\scriptstyle{\pm 0.028}}$
Table 5: Comparison on different checkpoints based on DDP-H5 methods.
A.2 Comparison on Full Demonstrations
To investigate how MaIL and Transformers scale with dataset size, we evaluated both models on the LIBERO benchmark with access to the full dataset. Due to time constraints, we focused solely on the Encoder-Decoder architecture for both MaIL and Transformers. We conducted experiments across different checkpoints, evaluating each model at three checkpoints and reporting their best performance. The results are presented in Table 6:
w/o language w/ language
Policy Backbone LIBERO-Object LIBERO-Spatial LIBERO-Long LIBERO-Goal
BC-H1 ED-Tr $0.847\scriptstyle{\pm 0.020}$ $0.658\scriptstyle{\pm 0.051}$ $\mathbf{0.658\scriptstyle{\pm 0.035}}$ $0.668\scriptstyle{\pm 0.055}$
ED-Ma $\mathbf{0.862\scriptstyle{\pm 0.036}}$ $\mathbf{0.693\scriptstyle{\pm 0.061}}$ $0.642\scriptstyle{\pm 0.011}$ $\mathbf{0.793\scriptstyle{\pm 0.050}}$
BC-H5 ED-Tr $\mathbf{0.910\scriptstyle{\pm 0.018}}$ $\mathbf{0.733\scriptstyle{\pm 0.049}}$ $0.632\scriptstyle{\pm 0.056}$ $0.772\scriptstyle{\pm 0.026}$
ED-Ma $0.907\scriptstyle{\pm 0.019}$ $0.697\scriptstyle{\pm 0.062}$ $\mathbf{0.687\scriptstyle{\pm 0.038}}$ $\mathbf{0.833\scriptstyle{\pm 0.015}}$
DDP-H1 ED-Tr $\mathbf{0.918\scriptstyle{\pm 0.028}}$ $0.707\scriptstyle{\pm 0.035}$ $0.683\scriptstyle{\pm 0.015}$ $0.788\scriptstyle{\pm 0.056}$
ED-Ma $0.913\scriptstyle{\pm 0.034}$ $\mathbf{0.725\scriptstyle{\pm 0.035}}$ $\mathbf{0.700\scriptstyle{\pm 0.015}}$ $\mathbf{0.818\scriptstyle{\pm 0.016}}$
DDP-H5 ED-Tr $\mathbf{0.952\scriptstyle{\pm 0.003}}$ $0.745\scriptstyle{\pm 0.018}$ $0.731\scriptstyle{\pm 0.031}$ $0.873\scriptstyle{\pm 0.030}$
ED-Ma $0.932\scriptstyle{\pm 0.029}$ $\mathbf{0.758\scriptstyle{\pm 0.045}}$ $\mathbf{0.772\scriptstyle{\pm 0.004}}$ $\mathbf{0.893\scriptstyle{\pm 0.034}}$
Table 6: Performance on LIBERO benchmark with access to the full dataset.
A.3 Experiments with Language Instructions
In Table 1 of our manuscript, we did not utilize language instructions for the LIBERO-Object, LIBERO-Spatial, and LIBERO-Long tasks. To gain further insights, we conducted additional experiments on these tasks using language embeddings, with access to 20% and 100% of the data. The results are presented in Tables 7 and 8:
w/ language
Policy Backbone LIBERO-Object LIBERO-Spatial LIBERO-Long
BC-H1 ED-Tr $0.596\scriptstyle{\pm 0.123}$ $0.474\scriptstyle{\pm 0.117}$ $0.316\scriptstyle{\pm 0.035}$
ED-Ma $\mathbf{0.689\scriptstyle{\pm 0.062}}$ $\mathbf{0.518\scriptstyle{\pm 0.080}}$ $\mathbf{0.408\scriptstyle{\pm 0.050}}$
BC-H5 ED-Tr $0.714\scriptstyle{\pm 0.036}$ $\mathbf{0.426\scriptstyle{\pm 0.035}}$ $0.278\scriptstyle{\pm 0.066}$
ED-Ma $\mathbf{0.764\scriptstyle{\pm 0.075}}$ $0.395\scriptstyle{\pm 0.083}$ $\mathbf{0.303\scriptstyle{\pm 0.115}}$
DDP-H1 ED-Tr $0.761\scriptstyle{\pm 0.005}$ $0.521\scriptstyle{\pm 0.002}$ $0.396\scriptstyle{\pm 0.031}$
ED-Ma $\mathbf{0.781\scriptstyle{\pm 0.045}}$ $\mathbf{0.576\scriptstyle{\pm 0.022}}$ $\mathbf{0.499\scriptstyle{\pm 0.048}}$
DDP-H5 ED-Tr $\mathbf{0.864\scriptstyle{\pm 0.025}}$ $\mathbf{0.618\scriptstyle{\pm 0.055}}$ $0.431\scriptstyle{\pm 0.010}$
ED-Ma $0.854\scriptstyle{\pm 0.040}$ $0.609\scriptstyle{\pm 0.050}$ $\mathbf{0.460\scriptstyle{\pm 0.035}}$
Table 7: Performance on LIBERO benchmark with 20% data.
w/ language
Policy Backbone LIBERO-Object LIBERO-Spatial LIBERO-Long
BC-H1 ED-Tr $0.803\scriptstyle{\pm 0.043}$ $0.718\scriptstyle{\pm 0.020}$ $0.724\scriptstyle{\pm 0.072}$
ED-Ma $\mathbf{0.859\scriptstyle{\pm 0.027}}$ $\mathbf{0.776\scriptstyle{\pm 0.075}}$ $\mathbf{0.770\scriptstyle{\pm 0.054}}$
BC-H5 ED-Tr $0.916\scriptstyle{\pm 0.027}$ $0.800\scriptstyle{\pm 0.018}$ $\mathbf{0.764\scriptstyle{\pm 0.039}}$
ED-Ma $\mathbf{0.929\scriptstyle{\pm 0.017}}$ $\mathbf{0.838\scriptstyle{\pm 0.048}}$ $0.758\scriptstyle{\pm 0.070}$
DDP-H1 ED-Tr $\mathbf{0.916\scriptstyle{\pm 0.025}}$ $\mathbf{0.813\scriptstyle{\pm 0.082}}$ $0.763\scriptstyle{\pm 0.037}$
ED-Ma $0.901\scriptstyle{\pm 0.020}$ $0.743\scriptstyle{\pm 0.043}$ $\mathbf{0.786\scriptstyle{\pm 0.030}}$
DDP-H5 ED-Tr $\mathbf{0.955\scriptstyle{\pm 0.005}}$ $0.836\scriptstyle{\pm 0.045}$ $0.826\scriptstyle{\pm 0.035}$
ED-Ma $0.943\scriptstyle{\pm 0.050}$ $\mathbf{0.883\scriptstyle{\pm 0.005}}$ $\mathbf{0.830\scriptstyle{\pm 0.063}}$
Table 8: Performance on LIBERO benchmark with full data.
Appendix B Additional Task Details
B.1 LIBERO benchmark
B.2 Real Robot Tasks
B.2.1 CupStacking
B.2.2 Inserting
B.2.3 PickPlacing
Appendix C Architectures
C.1 State Space Models
@@ Appendix E Real Robot Details
Figure 12: Data collection scene for real robot
@@ E.4 Evaluation
Appendix F Analysis on the State Space Representations
To better understand the advantages of MaIL over Transformers, we conducted a detailed analysis of the latent representations produced by both methods. Specifically, we compared the Encoder-Decoder architectures of MaIL and Transformers by visualizing their high-dimensional representations using t-SNE [46]. We employed BC-based models trained on the LIBERO-Object dataset using full demonstrations and rolled out the models across entire trajectories. The visualizations represent the latent spaces before the final action prediction layer (linear). Figure 13 displays the t-SNE results for four different trajectory representations.
As shown in the figure, MaIL’s state representations are notably more structured across all trajectories compared to those generated by Transformers. More structured representations could potentially improve the generalization of MaIL, which explains the advantages of MaIL over Transformers.
(a) Trajectory 1
(b) Trajectory 2
(c) Trajectory 3
(d) Trajectory 4
Figure 13: Visualization of latent representations on different trajectories.